OpenStudy (anonymous):

Which is not an equation of the line that passes through the points (1, 1) and (5, 5)? y - x = 0 x - y = 0 -x - y = 0 Which is not an equation of the line that passes through the points (1, 1) and (5, 5)? y - x = 0 x - y = 0 -x - y = 0 @Mathematics

OpenStudy (anonymous):

-x-y=0 is not an equation

OpenStudy (turingtest):

-x - y = 0 is the answer b/c -1-1=-2 not 0 -5-5=-10 not 0

OpenStudy (alfie):

I'd like to point out something. A line is a function you can invert on all R, since we know it to be bijective. Bijectivity means the function is also injective, so for each element of the domain, you have a DIFFERENT element of the codomain. This means, that if you give a x to your function, you can find just one y. You can now draw your conclusion on how to find it :)